# On the Robustness of Fully-Spiking Neural Networks in Open-World Scenarios using Forward-Only Learning Algorithms

### v1

@@ 2.4 Explainability and Interpretability
Contrasting with these approach, which rely solely on gradient information, Martinez et al. presented a technique involving a backward pass of a latent vector through the network to compute the relevance of specific features in the input space [24]. This method operates by reconstructing spike patterns across the layer and enabling the quantification of feature activity, thereby identifying regions of significance within the image. By doing so, the authors place additional focus on exploiting the representational capabilities of the latent space of non-convolutional networks to create interpretable techniques. Our paper follows the same geometrical motivation, leveraging the specific latent topology of FFA to provide enhanced interpretable results.
@@ 2.5 Contributions
This work presents a seamless adaptation of FFA for spiking models, offering an alternative formulation to the one proposed by Ororbia in [6]. In our formulation, we introduce two alternative goodness functions: an unbounded goodness function, which follows Ororbia’s methodology by treating spikes as a discrete version of ReLU activation; and a bounded goodness function, which fully leverages the bounded behavior of spiking models, making it more suitable for neuromorphic computers. Subsequently, we develop FF-SCP, an OoD detection algorithm that capitalizes on the characteristic latent space observed by the community to enhance the robustness of FFA. The applicability of FFA’s geometrical properties remains impactful for creating software that meets industry requirements in terms of security. However, the current trend in the literature has primarily focused on its predictive capabilities. Thus, this work offers an initial step toward providing additional methodologies that bring FFA closer to being a robust alternative to BP in terms of functionality. Additionally, to further explore the functional aspects of the latent space of FFA, we develop an attribution algorithm that provides understandable explanations for the features of an image that push it outside the distribution of any given class. This work also contributes to expanding the literature advocating for the study of the geometrical properties of latent spaces to benefit model reliability, particularly from the explainability field. However, we acknowledge the intrinsic challenge posed by the lack of better-suited architectures when introducing FFA into production-ready software.
@@ 3 Proposed Approach
This section introduces the methodological aspects of the algorithms proposed in this work, dedicating one subsection to each of the contributions. Subsection 3.1 outlines the adaptation of FFA for SNNs. Subsection 3.2 introduces the FF-SCP algorithm, which employs the latent state of SNNs trained with FFA to create an OoD detection algorithm. Lastly, Subsection 3.3 presents an attribution technique aimed at providing visual insights into the information relating arbitrary samples to the distribution of specific classes.
@@ 3.1 Spiking Forward-Forward Algorithm
Figure 1: Diagram illustrating the architecture, from left to right, of Backpropagation (BP), Feedback Alignment [31], and Forward-Forward Algorithm (FFA) [3]. Each architecture highlights the input forward path (black arrows), and the error propagation path (blue arrows). Additionally, we describe the update mechanism of FFA and its spiking adaptations.
Although the original motivation that led to the development of FFA was to propose a biologically plausible alternative to BP, the vast majority of research on this method is still obtained using ANNs. In this section, we present a straightforward adaptation of FFA for SNNs, modifying the goodness scoring function to aggregate the temporal dimension and employing surrogate gradient techniques to enable training under standard ML pipelines. Additionally, to capitalize on the bounded dynamics offered by spiking models, we propose a novel bounded goodness function that operates independently of the total simulation time. For the remainder of this paper, we will denote the spiking formulation of FFA as sFFA.
As detailed in Section 2, the original goodness function $G$ was formulated using the squared Euclidean distance of the latent vector. This function, combined with the unbounded dynamics of ReLU activations, enabled learning dynamics that could result in positive latent vectors achieving arbitrarily high goodness scores as training progressed. However, due to the discrete event-driven nature of SNNs, where the latent space consists of binary latent vectors, the activation value of each neuron is bounded by the total number of timesteps in the simulation. Therefore, to replicate behaviors similar to those observed in ANNs, the goodness function must be adapted to handle these discrete binary states. This paper introduces two implementations of the goodness function for spiking models: an unbounded goodness, which mimics the behavior of ANNs through the aggregation of the temporal dimension; and a bounded goodness, designed to capitalize on the bounded activation range of SNNs.
The first goodness function, namely the unbounded goodness function, referred to as $G_{\infty}(\cdot)$, can be naturally adapted from the conventional goodness function by considering the sum of spikes as a linear approximation of ReLU-like functions. In this scenario, the goodness function takes the latent vector $\bm{\ell}$, composed of the spike train of the neurons of the layer, and acts as the squared Euclidean norm of the number of spikes. Formally, given the latent vector $\bm{\ell}\in\{0,1\}^{N\times T}$ consisting of $N$ neurons and $T$ timesteps, the function $G_{\infty}(\ell):\{0,1\}^{N\times T}\rightarrow\mathbb{R}_{\geq 0}$ is formulated as:
@@
Our second goodness function, namely the bounded goodness, denoted by $G_{0}(\cdot)$, offers an alternative formulation in which the learning process relies solely on bounded states, aligning more closely with biological models compared to their unbounded counterparts. To achieve this, we replace the squared Euclidean distance with the Manhattan distance of the mean number of spikes in the latent spike train, ensuring all neurons contribute equally to the goodness score. Moreover, by applying mean aggregation, the obtained goodness scores remain independent of the total number of timesteps. Formally, given the latent vector $\bm{\ell}\in\{0,1\}^{N\times T}$, the bounded goodness function $G_{0}(\bm{\ell}):\{0,1\}^{N\times T}\rightarrow\mathbb{R}_{\geq 0}$ is expressed as:
@@
Upon early inspection of both goodness functions, several properties can be noted. On one hand, the unbounded function $G_{\infty}(x)$ is not strictly unbounded, as it spans an interval defined by the number of neurons $N$ and the total timesteps $T$, $[0,N\cdot T^{2}]$. However, the upper limit of this interval can achieve sufficiently high values to approximate the behavior of unbounded models. Moreover, this upper limit can be easily adjusted by changing the values of these two parameters, allowing for higher scores to approximate the behavior of unbounded models. On the other hand, the bounded design of $G_{0}(x)$ results in a range confined to the interval $[0,1]$, regardless of the aforementioned parameters. This bounded limit can pose challenges in achieving extreme values of the probability function $P(\cdot)$, especially if the original threshold $\theta$ was intended for unbounded goodness scores, as the function’s range will not properly span the probability range $[0,1]$. To address this, we modify the probability function in two ways. First, we introduce a scalar factor $\alpha$ to scale the term $G(\ell)-\theta$. Second, we replace the single threshold $\theta$ with two thresholds, $\theta_{\oplus}$ and $\theta_{\ominus}$, applying independent thresholds to the positive and negative probability functions, respectively. This adjustment allows us to increase $\alpha$ and adjust $\theta_{\oplus},\theta_{\ominus}$ to align with the domain of the probability function. Under this formulation, the resulting probability function is expressed as follows:
@@
3.2 The FF-SCP Algorithm
Figure 2: Visual diagram providing an example of the mechanics of our FF-SCP algorithm on an OoD sample using the latent space of classes $1$ and $3$. Additionally, we depict our interpretability method to obtain the attribution map of the OoD sample against the latent manifold of the class $1$.
This section introduces the FF-SCP algorithm, which leverages the emergent representational properties in the latent space of FFA to detect OoD inputs. As previously mentioned, multiple studies have observed that the latent space of networks trained using FFA exhibits a distinctive topology, characterized by distinct clusters for each training class [5, 4]. Conversely, due to ReLU activations having a minimum value of 00, negative data tends to generate near-zero latent vectors, resulting in all negative data clustering around the origin of the latent space. However, due to the inherent noise present in spiking models, negative latents attain higher activity bounds, shifting the single cluster of negative data into multiple different clusters dispersed throughout the space. These clusters retain the representability of the space, as each one can be observed to contain samples from the same class and which employed the same embedding label. Despite the presence of a high number of clusters, the neural specialization of the neurons contributes to creating a highly separated latent space, which is a desirable property for detecting samples not drawn from the training distribution. A visual representation of the emergent latent spaces is provided in Appendix B.
In its original formulation, the SCP algorithm proposed using a class-representative latent vector to describe each class, so that the geometric complexity of the latent space could be reduced into a single characteristic vector. To achieve this, SCP employed the centroid of the class latent vectors, allowing for straightforward computation of the OoD score by measuring the distance between the centroids and the sample’s latent vector representation. However, this representative vector failed to perform adequately when presented with data distributions whose latent subspaces resulted in separated topologies or highly unbalanced latent vector distributions, as the center of the data did not capture the complex geometry of these latent manifolds.
@@ 3.2 The FF-SCP Algorithm
To mitigate this effect, this work proposes an alternative distance heuristic, replacing the centroid distance with a point-set distance between the latent vector $\bm{\ell}\in\mathbb{R}^{N}$ of a sample $x$ and a set of latent vectors $\mathcal{L}_{c}=\{\bm{\ell}^{1},\bm{\ell}^{2},\dots\}$ extracted from the complete latent manifold of class $c$, where $N$ is the dimension of the layer. This distance function is formulated using a prior distance function $d(\bm{z}_{1},\bm{z}_{2})$, which measures the distance between two arbitrary vectors $z_{1},z_{2}\in\mathbb{R}^{n}$. The distance between a point and a set is then defined as the minimum distance between the point and any point in the set. From a theoretical perspective, this function should operate over the complete latent manifold $\overline{\mathcal{L}_{c}}\subset\mathbb{R}^{N}$, however, due to the discrete nature of current AI, these clusters can only consist of a finitely many set of points. Therefore, we approximate this distance using the following expression:
$d(\bm{\ell},\overline{\mathcal{L}_{c}})=\inf_{\bm{\ell}\in\overline{\mathcal{L}}_{c}}d(\bm{\ell},\bm{\ell})\approx\min_{\bm{\ell}\in\mathcal{L}_{c}}d(\bm{\ell},\bm{\ell}).$ (7)
@@
Given that FFA provides different representations of the same input sample $x$ depending on the embedded label $p$ used during the supervised inference process, the input data used to create the latent space representatives must include both positive and negative instances of input data. Therefore, creating the set $\mathcal{L}$ of the different clusters must involve a uniform amount of inputs from each class using all possible labels and storing them. Once all these vectors have been captured, the set $\mathcal{L}$ will provide a sufficient representation of the original latent space. [^1] Given that networks trained with FFA do not currently achieve state-of-the-art accuracy, a filtering step is necessary to remove potentially misclassified latent vectors. This step involves discarding a small set of the least active elements from sets where the embedded label aligns with the real class, as they may have been misidentified as negative samples. Conversely, it also involves removing a reduced set of the latent vectors with the highest activity from the negative sample set to avoid potential misclassification of these highly active elements as positive samples. To account for the presence of negative vectors, we extend the definition of $\mathcal{L}$ to $\{\mathcal{L}_{c,p}\mid c\in\mathcal{C}\text{ and }p\in\mathcal{C}\}$, where $\mathcal{L}_{c,p}$ represents the set of latent vectors from class $c$ forwarded using the embedded label $p$, and $\mathcal{C}$ denotes the set of class labels.
@@
Once the representative set of latent vectors has been identified, an initial version of the scoring function can be formulated as the distance between the sample’s latent vector and each class-representative set of clusters. Unlike conventional neural networks, where the latent representation of a sample is unique, FFA generates different representations depending on the embedded label used during the forward process. Therefore, to measure the distance between a sample and the set of class-representative clusters, it’s crucial to account for the different representations that the sample obtains with different labels. Thus, the distance from the latent vector to each class is defined as the sum of the distances from the vector to each of the class clusters. This effect is illustrated in Figure 2, where the red lines define the distances from each latent vector of the different embeddable labels with respect to the latent representative sets $\mathcal{L}_{c,p}$.
@@
Consequently, the score of the unknown sample $x$ is defined as the minimum distance from its latent representation to each class representative set of clusters. Once this minimal distance is computed, ID samples are expected to achieve a low overall distance, as their representation should remain proximal to their respective class set of clusters. Conversely, OoD samples are expected to attain a higher minimal distance, as their latent vectors should not be within proximity to any set of ID clusters. Therefore, the initial score function can be defined as:
$s(x)=\min_{p\in\mathcal{C}}\sum_{c\in\mathcal{C}}d\left(\mathcal{N}_{d}(x,c),\mathcal{L}_{p,c}\right)^{\beta},$ (8)
@@
where $\mathcal{N}_{d}(x,p)$ represents the output latent vector obtained from the forward pass of an input $x$ with the embedded label $p$ at depth $d$, and $\beta$ denotes a scaling hyperparameter.
When working with certain spaces, this scoring function can encounter a significant limitation, as an insufficient number of samples can result in inaccuracies when measuring the discrete approximation of the distance. For instance, in scenarios where the latent representation contains widely dispersed points, the resulting measurements will tend to lean towards large distances, as the distance approximation might fail to find nearby neighbors. This issue is more pronounced in high-dimensional spaces, where the curse of dimensionality requires exponentially more samples to achieve even distributions of points as the dimensions increase. Furthermore, due to the activity maximization objective of FFA, where positive latent vectors strive to maximize their distance from the origin, class manifolds can see their diameters extended over large distances, making it harder to achieve accurate representations.
@@
In such scenarios, an extreme outcome can occur where the scoring of samples is inverted, leading to OoD samples achieving lower scores than ID samples. This situation arises when the minimal distance from the zero vector to any representative cluster is smaller than the distance between points within the same representative cluster. Consequently, OoD samples, typically characterized by nearly zero latent activity, can end up with lower scores than ID samples with high latent activity, simply because they are unable to find a nearby neighbor within their own cluster to minimize the scoring function. Formally, this issue arises when the following inequality holds:
$\min_{c\in C}d(\textbf{0},L_{c,c})\leq\mathbb{E}\left[\min_{c\in C}d(\mathcal{N}_{p}(\bm{X},c),\mathcal{L}_{c,c})\right],$ (9)
@@
where $\bm{X}$ represents a randomly selected sample in the ID dataset. A more detailed explanation of this result is presented in Appendix A.
@@
To mitigate this issue, an additional branch is incorporated into the function, so that the samples satisfying this condition are handled separately. Within this branch, instances closer to the zero vector than to any other class cluster are assigned a large value $\gamma$. This designation indicates that these instances are more likely to belong to the negative distribution, which consists of synthetic samples outside the training distribution. However, we acknowledge that a small proportion of ID samples may also fall into this category. For these instances, a greater distance from the zero vector suggests a higher likelihood of misclassification. Therefore, we introduce a negative $s(x)$ term into the expression so that elements that fall into this branch have their order reversed.
@@
Therefore, after the proposed changes, the final scoring function is given by:
@@
where $\gamma$ denotes a constant value designed to assign an elevated score to samples falling into the first branch, the term $s_{0}(x)$ represents is equivalent to the function $s(\cdot)$ but replacing the spaces $\mathcal{L}_{i,j}$ for the space comprised by the vector $\bm{0}$, and $z_{s}$ serves as a scaling factor for the zero distance $s_{0}$.
Once the scores of the data are computed, the labeling of OoD samples follows a conventional process. This is achieved by comparing the score obtained by a sample with a manually chosen threshold value. If the score of the sample is greater than the threshold value, the sample is classified as OoD; conversely, if the score is less than the threshold, the sample is identified as ID. The threshold value is user-defined, depending on the desired confidence level for the model to operate. Therefore, the final heuristic used with FF-SCP is detailed in Algorithm 1.
1
2
Data: Batch of input values $X$ and their respective labels $Y$.
Result: Distances to the set of averaged feature vectors
3
4
5Function initialize_latents($X$, $Y$):
       $\triangleright$ Latent Representative Generation
6       for each $(x,y)$ in $(X,Y)$ do
7             for each $p$ in $\mathcal{C}$ do
8                   Obtain the latent vector $\ell$ from $\mathcal{N}_{1}$(x, p)
9                   Append $\ell$ into the set $L_{y,p}$
10
11
       $\triangleright$ Filtering Step
12       for each $(p,c)$ in $\mathcal{C}\times\mathcal{C}$  do
13             if p equals c then
14                   Remove the lowest normed elements of $\mathcal{L}_{p,c}$
15
16            else
17                   Remove the highest normed elements of $\mathcal{L}_{p,c}$
18
19
20
21
22 Function get_score($x$):
23
24      Define the distance vector $\bm{D}$ as $|C|$ zeros
25       Define the zero distance scalar $D^{0}$ to zero
26
27      for each $c$ in $\mathcal{C}$ do
28             Get the latent vector $\bm{\ell}$ from $\mathcal{N}_{1}$($x$, $c$)
29
30            Add the value $d(\bm{\ell},\bm{0})$ to $D^{0}$
31
32            for each $p$ in $C$ do
33                   Add $d(\bm{\ell},\mathcal{L}_{c,p})$ to the distance to the class $p$, $D_{p}$
34
35
36      Let the score $s$ be the minimal value in the vector $\bm{D}$
37       return $s$ if $s\geq D^{0}$ else $\gamma-s$
38
39
Algorithm 1 FF-SCP

### v2

@@ 2.4 Interpretability
In their work, Martinez-Seras et al. (2023) present a technique suitable for spiking systems, involving a backward pass of a latent vector through the network to compute the relevance of specific features in the input space. This method operates by reconstructing spike patterns across the layers and enabling the quantification of feature activity, thereby identifying regions of significance within the image. The authors highlighted the importance of exploiting the representational capabilities of the latent space of non-convolutional networks to create interpretable techniques.
@@ 2.5 Contributions
This work introduces the FF-OoD algorithm, a spiking OoD detection method that leverages the characteristic latent topology observed in networks trained with FFA. While most existing literature on FFA has focused on its predictive capabilities in analog settings, this work takes an initial step toward exploring its applicability, paving the way for neuromorphic systems. Additionally, we propose an attribution technique that generates interpretable images, highlighting the features that push an image outside the distribution of any selected class. Through these developments, we aim to bridge the gap in the literature on OoD detection and interpretability in SNNs.
@@ 3 Proposed Approach
This section introduces the methodological details of the algorithms proposed in this work. Given the differences from current approaches, Section 3.1 outlines the adaptations employed in FFA to make it suitable for SNNs. Subsequently, Section 3.2 introduces the FF-OoD algorithm, which relies on the latent geometry of networks trained with FFA to create an OoD detection algorithm. Lastly, Section 3.3 presents an attribution technique aimed at providing visual insights into the information linking arbitrary samples to the distribution of specific classes.
@@ 3.1 Spiking Forward-Forward Algorithm
Figure 1: Diagram illustrating the architecture, from left to right, of Backpropagation (BP), Feedback Alignment (Lillicrap et al., 2016), and Forward-Forward Algorithm (FFA) (Hinton, 2022). Each architecture highlights the input forward path (black arrows), and the error propagation path (blue arrows). Additionally, we describe the update mechanism of FFA and its spiking adaptations.
Despite its original design as a biologically plausible alternative to Backpropagation, FFA has been mainly employed to train traditional ANNs. As previously mentioned, efforts have already been made to develop more biologically accurate alternatives, replacing its original error-based formula with a Hebbian Learning Rule. However, to avoid additional complexity within the network architecture, we advocate for the use of a surrogate gradient for the local gradient computation, which allows SNNs to be trained using the original learning heuristic. Following this design choice, the rest of the subsection details the modifications made to the goodness functions to adapt them to the new time dimension.
As detailed in Section 2, the original goodness function $G$ was formulated using the squared Euclidean distance of the latent vector. This function, combined with the unbounded dynamics of ReLU activations, enabled learning dynamics that could result in positive latent vectors achieving arbitrarily high goodness scores as training progressed. However, due to the discrete event-driven nature of SNNs, where the latent space consists of binary latent vectors, the activation value of each neuron is bounded by the total number of achievable spikes during the simulation period. Therefore, to replicate behaviors similar to those observed in ANNs, the goodness function must be adapted to handle these discrete states. We introduce two goodness functions for spiking models: an unbounded goodness, which mimics the behavior of ANNs through the aggregation of the temporal dimension; and a bounded goodness, designed to capitalize on the bounded activation range of SNNs.
Unbounded Goodness
This goodness function, referred to as $G_{\infty}(\cdot)$, is obtained from the conventional goodness function by aggregating the spikes, resulting in an approximation of a ReLU-like function. In this scenario, the goodness function takes the latent vector $\bm{\ell}$, composed of the spike train of the neurons of the layer, and acts as the squared Euclidean norm of the number of spikes. Formally, given the latent vector $\bm{\ell}\in\{0,1\}^{N\times T}$ consisting of $N$ neurons and $T$ timesteps, the function $G_{\infty}(\ell):\{0,1\}^{N\times T}\rightarrow\mathbb{R}_{\geq 0}$ is formulated as:
@@
Although this function is not strictly unbounded, as it spans the interval $[0,N\cdot T^{2}]$, the upper limit of this interval can be reached by setting sufficiently high values of $N$ and $T$ to mimic the behavior of unbounded models.
Bounded Goodness
Our second goodness function, namely the bounded goodness, denoted by $G_{0}(\cdot)$, employs the average spiking rates of the neurons within the layer. By using this aggregation method, the obtained score remains independent of the number of timesteps of the simulation. Using the previous notation, the function $G_{0}(\bm{\ell}):\{0,1\}^{N\times T}\rightarrow\mathbb{R}_{\geq 0}$ is expressed as:
@@
Naturally, the bounded design of $G_{0}(x)$ operates in a range limited to $[0,1]$, regardless of the hyperparameter configuration. This bounded limit can pose a challenge in achieving extreme values for the probability function $P(\cdot)$, especially if the original threshold $\theta$ was intended for unbounded goodness scores. To address this, we modify the probability function in two ways. First, we introduce a scalar factor $\alpha$ to scale the term $G(\ell)-\theta$. Second, we replace the single threshold $\theta$ with two thresholds, $\theta_{\oplus}$ and $\theta_{\ominus}$, applying independent thresholds to the positive and negative probability functions, respectively. This adjustment allows us to increase $\alpha$ and adjust $\theta_{\oplus},\theta_{\ominus}$ to align with the domain of the probability function. Under this formulation, the resulting probability function is expressed as follows:
@@
3.2 FF-OoD for Out-of-Distribution Detection
Figure 2: Visual diagram providing an example of a) the mechanics FF-OoD on an ID sample (green) and a OoD sample (red), with respect to the set of data from classes $1$ and $3$, and 2) the process of optimizing a latent vector (purple) so that when decoded, it can be compared with the original sample (the OoD sample) to obtain the atribution map to some specified class (Class 1).
As shown in multiple works (Ororbia and Mali, 2023; Tosato et al., 2023), one key feature observed in networks trained using FFA is its emerging latent topology. This space has been characterized by a set of highly separated latent clusters, each related to a different class. This property, together with the observed sparsity of the latent vectors, showcases the discriminative capacity of the neurons within each layer of the network. In this section, we exploit this geometrical landscape in order to develop a OoD scoring function.
An additional property of FFA observed in our work is that class clustering is not solely confined to the positive latent space, but also appears in networks trained with bounded activations. Due to its original formulation, the negative goodness minimization process combined with ReLU activations resulted in the negative latent space being compressed into a small cluster around the zero vector. However, in scenarios where the inputs are noisy or when activations are bounded, the negative data exhibit non-zero latent vectors, which form their own clusters. These clusters can be characterized by the pair consisting of the real class label and the negative label embedded within the input. A visual representation of the emergent latent spaces is provided in C.
Building on the geometrical motivation of SCP, we develop a distance-based algorithm that utilizes networks trained with FFA. In their work, Martinez-Seras et al. relied on the distance from an input latent vector to the set of class representatives, each defined by the centroid of the precomputed class latent distribution. However, reducing the geometry of a latent space to its centroid can be too restrictive when working with non-trivial connected latent manifolds or separated manifolds. In such cases, a single point may not serve as an adequate measure of distance from an arbitrary point in the space with respect to the class distribution. For example, when the manifolds are not connected, these centroids often lie outside the respective parts of the space.
@@ 3.2 FF-OoD for Out-of-Distribution Detection
To mitigate this effect, this work replaces the distance to the centroid with a point-set distance between the latent vector $\bm{\ell}\in\mathbb{R}^{N}$ of a sample $x$ and a set of latent vectors $\mathcal{L}_{c}=\{\bm{\ell}^{1},\bm{\ell}^{2},\dots\}$ extracted from the complete latent manifold of class $c$, where $N$ is the dimension of the layer. This distance function is formulated using a prior distance function $d(\bm{z}_{1},\bm{z}_{2})$, which measures the distance between two arbitrary vectors $z_{1},z_{2}\in\mathbb{R}^{n}$. The distance between a point and a set is then defined as the minimum distance between the point and any point in the set. From a theoretical perspective, this function should operate over the complete latent manifold $\overline{\mathcal{L}_{c}}\subset\mathbb{R}^{N}$, however, due to the discrete nature of current AI, these clusters can only consist of a finite set of points. Therefore, we define the empirical distance $\widehat{d}(\cdot)$ by approximating the real manifold following the expression:
$d(\bm{\ell},\overline{\mathcal{L}_{c}})=\inf_{\bm{\ell}_{t}\in\overline{\mathcal{L}}_{c}}d(\bm{\ell},\bm{\ell}_{t})\approx\min_{\bm{\ell}_{t}\in\mathcal{L}_{c}}d(\bm{\ell},\bm{\ell}_{t})=\widehat{d}(\bm{\ell},\mathcal{L}_{c}).$ (7)
@@
Given that FFA provides different representations of the same input sample $x$ depending on the embedded label $p$ used during the supervised inference process, the input data used to create the latent space representatives must include both positive and negative instances of input data. Therefore, creating the set $\mathcal{L}$ of the different clusters must involve a uniform amount of inputs from each class, using all possible labels. Once all these vectors have been captured, the set $\mathcal{L}$ will provide a sufficient representation of the original latent space. Nevertheless, given that networks trained with FFA do not currently achieve state-of-the-art accuracy, a filtering step is necessary to remove potentially misclassified latent vectors. This step involves discarding a small number of the least active elements from sets where the embedded label aligns with the real class, as they may have been misidentified as negative samples. Conversely, it also involves removing a reduced set of the latent vectors with the highest activity from the negative sample set to avoid potential misclassification of these highly active elements as positive samples. To account for the presence of negative vectors, we extend the previous definition of $\mathcal{L}$ for $\{\mathcal{L}_{c,p}\mid c\in\mathcal{C}\text{ and }p\in\mathcal{C}\}$, where $\mathcal{L}_{c,p}$ represents the set of latent vectors from class $c$ forwarded using the embedded label $p$, and $\mathcal{C}$ denotes the set of class labels.
@@
Once the representative set of latent vectors has been identified, an initial version of the scoring function can be formulated as the distance between the sample’s latent vector and each class-representative set of clusters. Unlike conventional neural networks, where the latent representation of a sample is unique, FFA generates different representations depending on the embedded label used during the forward process. Therefore, to measure the distance between a sample and the set of class-representative clusters, it’s crucial to account for the different representations that the sample obtains with different labels. Thus, the distance from the latent vector to each class is defined as the sum of the distances from the vector to each of the class clusters. This effect is illustrated in Figure 2, where the green lines of the ID sample define the distance from each latent vector of the different embeddable labels with respect to the latent representative sets $\mathcal{L}_{c,p}$.
@@
Consequently, the score of the unknown sample $x$is defined as the minimum distance between its latent representation, using the different possible input labels, and each set of class-representative clusters. Once this distance is computed, ID samples are expected to achieve a low overall distance, as their representation should remain close to their respective class set of clusters. In contrast, OoD samples are expected to attain a higher minimal distance, as their latent vectors should not be close to any set of ID clusters. Thus, the initial score function can be defined as
$s(x)=\min_{p\in\mathcal{C}}\sum_{c\in\mathcal{C}}\widehat{d}\left(\mathcal{N}_{k}(x,c),\mathcal{L}_{p,c}\right)^{\beta},$ (8)
@@
where $\mathcal{N}_{k}(x,p)$ represents the output latent vector obtained from the forward pass of an input $x$ with the embedded label $p$ at depth $k$, and $\beta$ denotes a scaling hyperparameter.
When working with certain spaces, this scoring function can encounter significant limitations, as an insufficient number of samples can lead to inaccuracies in the discrete approximation of the distance. For instance, in scenarios where the latent representation covers a large volume with low density, the large distance between points may cause the distance function to fail in finding any close neighbors, resulting in a high distance value. This issue is especially pronounced in high-dimensional spaces due to the curse of dimensionality, as evenly distributed representations require an exponential number of samples.
@@
In such scenarios, an extreme outcome can occur where the scoring of samples is inverted, leading to OoD samples achieving lower scores than ID samples. This situation arises when the minimum distance from the zero vector to any representative cluster is smaller than the distance between points within the same representative cluster. Consequently, OoD samples, typically characterized by nearly zero latent activity, can end up with lower scores than ID samples with high latent activity, simply because they are unable to find a nearby neighbor within their own cluster to minimize the scoring function. Formally, this issue arises when the following inequality holds:
$\min_{c\in C}\widehat{d}(\textbf{0},L_{c,c})\leq\mathbb{E}\left[\min_{c\in C}\widehat{d}(\mathcal{N}_{p}(\bm{X},c),\mathcal{L}_{c,c})\right],$ (9)
@@
where $\bm{X}$ represents a randomly selected sample in the ID dataset. A more detailed explanation of this result is presented in A.
@@
To mitigate this issue, an additional branch is incorporated into the original function to handle samples that meet this condition separately. Within this branch, instances closer to the zero vector than to any other class cluster are assigned a large value $\gamma$. This high value indicates that these instances are more likely to belong to the negative distribution, which consists of synthetic samples outside the training distribution. However, we acknowledge that a small proportion of ID samples may also fall into this category. For these instances, a greater distance from the zero vector suggests a higher likelihood of misclassification. Therefore, we introduce a negative $s(x)$ term into the expression, ensuring that elements that fall into this branch have their order reversed.
@@
Therefore, after the proposed changes, the final scoring function is defined by:
@@
where $\gamma$ denotes a constant value designed to assign an elevated score to samples falling into the first branch, the term $s_{0}(x)$ represents is equivalent to the function $s(\cdot)$ but replacing the spaces $\mathcal{L}_{i,j}$ for the space comprised by the vector $\bm{0}$, and $z_{s}$ serves as a scaling factor for the zero distance $s_{0}$. This final heuristic used with FF-OoD is detailed in Algorithm 1.
Algorithm 1 FF-OoD
1:  INITIALIZE_LATENTS($X$, $Y$):
2:        for each $(x,y)$ in $(X,Y)$:
3:            for each $p$ in $\mathcal{C}$:
4:                Obtain the latent vector $\ell$ from $\mathcal{N}_{1}(x,p)$
5:                Append $\ell$ into the set $\mathcal{L}_{y,p}$
6:        for $(p,c)$ in $\mathcal{C}\times\mathcal{C}$
7:            if $p=c$
8:                Remove samples of $\mathcal{L}_{p,c}$ with low norm
9:            else
10:                Remove samples of $\mathcal{L}_{p,c}$ with high norm
11:  GET_SCORE($x$):
12:        Define the distance vector $\bm{D}$ as $|\mathcal{C}|$ zeros
13:        Define the zero distance scalar $D^{0}$ to zero
14:        for each $c$ in $\mathcal{C}$:
15:            Get the latent vector $\bm{\ell}$ from $\mathcal{N}_{1}(x,c)$
16:            Add the value $d(\bm{\ell},\bm{0})$ to $D^{0}$
17:            for $p$ in $\mathcal{C}$:
18:                Add $\widehat{d}(\bm{\ell},\mathcal{L}_{c,p})$ to $D_{p}$
19:        Let the score $s$ be the minimal value in the vector $\bm{D}$
20:        return $s$ if $s\geq D^{0}$ else $\gamma-s$